OpenStudy (anonymous):

please helpp mee ): In a geometric sequence, if r=square root of 3 and a4=2,what is the value of a1?

OpenStudy (anonymous):

\(a_1,a_1r,a_1r^2,a_1r^3\) \[a_1r^3=a_4\] \[a_1\times 3\sqrt{3}=a_4=2\]

OpenStudy (anonymous):

so \[a_1=\frac{2}{3\sqrt3}\]

OpenStudy (anonymous):

it says the answer is 3radical3

OpenStudy (anonymous):

):

OpenStudy (anonymous):

Check your question and your approved solution again. Pretty sure satellite73 is spot on. The correct answer might have be simplified to 2sqrt3/9